OpenStudy (anonymous):

3^√x-3 – 5 = 16? plz exlpain how to solve.

OpenStudy (anonymous):

Is this: \[ \large f(x)=3^{\sqrt{x-3}}-5=16\]

OpenStudy (anonymous):

Iff write the equation like this: \[ \LARGE 3^{\sqrt{x-3}} = 21\] then use logarithm properties to solve for x.